A glass column is filled with mercury and inverted in a pool of mercury. The mercury column stabilizes at a height of 728 mm above the pool of mercury. What is the pressure of the atmosphere?

0.639 atm

0.958 atm

455 atm

1.04 atm

0.728 atm

A gas sample is held at constant pressure. The gas occupies 3.62 L of volume when the temperature is 21.6°C. Determine the temperature at which the volume of the gas is 3.37 L.

20.1 K

295 K

548 K

274 K

317 K

Gaseous chlorine is held in two separate containers at identical temperature and pressure. The volume of container 1 is 1.30 L, and it contains 6.70 mol of the gas. The volume of container 2 is 2.08 L. How many moles of the gas are in container 2?
Answer #1

1) The height of the mercury column = 728 mm

1 atm of pressure = 760 mm of the mercury column.

Hence,

728 mm = 728 / 760 mm

= 0.958 atm  

2) We will apply Charles law here

Vi / Ti = Vf / Tf ------------(i)

Ti = 21.6o C = (21.6 + 273 ) = 294.6 K

Substituting the values in equation i , we get

3.62/ 294.6 = 3.37 / Tf

=> Tf = 3.37 *294.6 / 3.62 = 274 K